i Data Table (In millions) | Total current assets $ ................ 15.5 15.8 Property, plant, equipment, and other.assets .a. Earned revenue, $2.5 million, on account. Now, include the a. transaction for 2019 in our original ratio calculations at Dd. Paid selling expense of $0.5 million. Now, include the d. transaction for 2019 in our original ratio calculations at DecemRequirement 3. Complete the statements with either increase or decrease. a. Revenues usually the current ratio. b. Revenu
Answer #1
a. Current Ratio = Current Assets/Current Liabilities
Current Ratio = (15.5+2.5)/9.8 = 1.84 times
Note : Revenue of 2.5 Million $ on account increases our trade receivables and hence current assets are increased by 2.5 Million $.
Debt ratio = Total debt/Total Assets
Debt ratio = (Current Liabilities + Long-term liabilities) / (Current Assets + Property Plant & Equipment)
Debt ratio = (9.8+5.3)/(15.5+2.5+15.8) = 44.67%
b. Current Ratio = Current Assets/Current Liabilities
Current Ratio = (15.5+6)/9.8 = 2.19 times
Note : Since we have borrowed 6 Million$ long-term debt, our cash balance also increases which increases our current assets as well.
Debt ratio = Total debt/Total Assets
Debt ratio = (Current Liabilities + Long-term liabilities) / (Current Assets + Property Plant & Equipment)
Debt ratio = (9.8+5.3+6)/(15.5+6+15.8) = 56.57%
c. Current Ratio = Current Assets/Current Liabilities
Current Ratio = (15.5-4.9)/(9.8-4.9) = 2.16 times
Note : Half of CL paid in cash = 9.8/2 = 4.9 Million $. It means both our current liabilities as well as current assets will go down by 4.9 Million $
Debt ratio = Total debt/Total Assets
Debt ratio = (Current Liabilities + Long-term liabilities) / (Current Assets + Property Plant & Equipment)
Debt ratio = (9.8-4.9+5.3)/(15.5-4.9+15.8) = 38.64%
d. Current Ratio = Current Assets/Current Liabilities
Current Ratio = (15.5-0.5)/9.8 = 1.53 times
Note : Selling expenses paid of 0.5 Million $ reduces cash and hence current assets by that amount.
Debt ratio = Total debt/Total Assets
Debt ratio = (Current Liabilities + Long-term liabilities) / (Current Assets + Property Plant & Equipment)
Debt ratio = (9.8+5.3)/(15.5-0.5+15.8) = 49.03%